I,m having trouble coming up with an equation to solve this problem. Could you please help?
At 7:18 you get on the bus and are off to school. The bus travels at 42mph. Fifteen minutes later your mother realized you left your math project on the kitchen table. She drives 72mph to catch up with the bus. What time will your mom catch up with the bus?Will she get you your project before the homeroom bell? FYI the homeroom bell in my school rings at 7:37.

Answer by Paul(988)

15 minutes = 0.25h
42(x+0.25)=72x
42x+10.5=72x
30x=10.5
x=0.35

0.35*60=21 minutes =7:39

Hence, she will reach there in 21 minutes and no she won't be able to reach on time.
Paul.